# Bribe & Fork: Cheap PCN Bribing Attacks via Forking Threat

### v1

@@ Abstract
In this work, we reexamine the vulnerability of Payment Channel Networks (PCNs) to bribing attacks, where an adversary incentivizes blockchain miners to deliberately ignore a specific transaction to undermine the punishment mechanism of PCNs. While previous studies have posited a prohibitive cost for such attacks, we show that this cost may be dramatically reduced (to approximately \$125), thereby increasing the likelihood of these attacks. To this end, we introduce Bribe & Fork, a modified bribing attack that leverages the threat of a so-called feather fork which we analyze with a novel formal model for the mining game with forking. We empirically analyze historical data of some real-world blockchain implementations to evaluate the scale of this cost reduction. Our findings shed more light on the potential vulnerability of PCNs and highlight the need for robust solutions.
@@ 1 Introduction
The financial world was transformed by blockchains such as Bitcoin [27] and Ethereum [35]. While blockchains offer a number of benefits, their scalability remains a significant challenge when compared to traditional centralized payment systems [15]. One promising solution to this issue is the so-called payment channel networks (PCNs) that move most of the transaction workload off-chain [20].
Several PCN proposals [6, 7, 8, 9, 11, 16, 17, 18, 19, 21, 28, 31] have been laid forward so far, each design offering some unique combination of features. Nonetheless, the core idea behind payment channels remains the same, that is to facilitate off-chain transactions among parties connected, either directly or indirectly via a path, on a network operating on top of the blockchain layer, the PCN.
@@
Hence, the security of PCNs, like the most widely deployed Bitcoin Lighting Network [28], crucially relies on financial incentives. Specifically, during the dispute period, the punishment mechanism should enforce that a malicious party is always penalized and an honest party should never lose its funds. Unfortunately, this is not always the case, as argued by Khabbazian et al. [26] and Avarikioti et al. [10]. For instance, Lighting channels are susceptible to the so-called timelock bribing attacks. In such an attack, a malicious party posts an old update transaction on-chain, attempting to close its channel with more funds than it presently possesses. Concurrently, the party bribes the miners to ignore the punishment transaction of the counterparty for the duration of the dispute period. This bribe is typically offered to the miners in the form of high transaction fees.
@@
Naturally, the success of this attack depends on the value of the bribe. Avarikioti et al. [10] showed that a bribing attack will be successful if the bribe is no smaller than: $\dfrac{f_{1}-f}{\lambda_{min}},$ where $\lambda_{min}$ is the fraction of the mining power controlled by the least significant miner in the underlying proof-of-work blockchain, $f_{1}$ is the sum of the fees of a block containing the punishment transaction, and $f$ is the sum of the fees of a block containing average transactions. We observe however that as $\lambda_{min}$ may be arbitrarily small, the bribing amounts required may significantly exceed the funds typically locked in PCNs, rendering the bribing attack impractical. Moreover, even in blockchains with rather concentrated mining power like in Bitcoin [22], the cost of a bribing attack is very high. For example, conservatively assuming that the smallest miner has $10^{-4}$ of the total mining power [^1] In the Section 0.D.2, we experimentally show, that the mining power of the weakest miner in the system can be fairly assumed to be of magnitude $10^{-12}$., the cost of the attack as analyzed in [10] would be at least $1$ BTC, for $f_{1}-f\approx 10^{-4}$ BTC. As a result, it would be irrational to perform a bribing attack of this sort, as the average closing price for $1$ BTC between, for instance, 2019 and 2022 was $23,530.92$ USD222statmuse.com, which is more than 10-fold the current total value locked on average in a Lighting channel [^3] https://1ml.com/statistics. This naturally leads to questioning whether there is potential to amplify such attacks to the extent they pose a genuine threat to the security of PCNs.
1.0.1 Our Contribution.
In this work, we show that bribing costs can be significantly reduced, thereby making timelock bribing attacks a realistic threat. We do so by extending the bribing attack to leverage not only the structure of transaction fees but also a threat to fork the blockchain, known as a feather fork attack [23, 30]. In our context, a given miner executes a feather fork attack by announcing a self-penalty transaction $tx_{p}$. Whenever the self-penalty transaction appears on the blockchain, the miner is incentivized to fork the punishment transaction $tx_{1}$ on the blockchain, i.e., the miner will try to extend the blockchain based on the predecessor of the block $tx_{1}$ including some other block. Specifically, a feather-forking miner is bribed to commit collateral, betting that their fork will win the race. As a consequence, their threat of forking becomes considerably credible, incentivizing other miners to follow their fork. The collateral is of a similar magnitude as the bribe in [10], however, the miners only lock it temporarily. We call our attack Bribe & Fork. With the feather fork at hand, the bribing cost may now be reduced from $\frac{f_{1}-f}{\lambda_{min}}$ to approximately:
@@
where $\overline{f}$ is the average fee of a single transaction, and $\lambda_{s}$ is the mining fraction of the most significant miner. Recall that $f_{1}$ and $f$ denote the sum of fees of a block containing the punishment transaction and only average transactions respectively. To demonstrate the cost reduction of Bribe & Fork, we reexamine the previous example for Bitcoin, with $\overline{f}\approx 10^{-4}$ BTC, $f_{1}-f\approx\overline{f}$, and $\lambda_{s}\approx 20\%$. Now, $\lambda_{s}$ replaces in the denominator the previously presented $\lambda_{min}<<10^{-4}$, thus yielding a bribe at least 1000 times smaller than the one in [10].
To derive this result, we present a formal model of mining games with forking, extending the conditionally timelocked games introduced in [10]. In the game with forks, miners may now choose, in each round, (a) which transactions to mine, and (b) whether they want to continue one of the existing chains or they intend to fork one of the chains. All miners know the choices of the winner of each block, as a feather-forking miner locks collateral on-chain.
To empirically estimate the cost reduction of Bribe & Fork, we analyze the historical data of real-world blockchain implementations. Among others, we analyze the average block rewards and fees, as well as the hash power present in the system and available to a single miner, primarily for Bitcoin in 2022. Given the officially available data, we observe that the cost of our attack can be as cheap as \$125 (for 1 BTC $\approx$ \$25.000). In general, the cost of our attack can be up to $10^{10}$ times cheaper than the bribe required in [10] according to our findings. Hence, even considering a collateral of around \$$\$30,000$, Bribe & Fork is substantially more cost-efficient, and by extension, more probable to occur.
2 The Bribe & Fork Attack
In this section, we first describe the necessary components of Bribe & Fork, including feather fork attacks as well as timelock bribing attacks in the setting of PCNs, and conclude with the description of the Bribe & Fork attack.
2.1 Necessary Components and Background
@@
Timelock Bribing in the Bitcoin Lightning Network. Next, we describe the timelock bribing attack in the context of the most widely deployed payment channel network, namely the Bitcoin Lightning Network (LN). In LN, a single on-chain transaction called the funding transaction, opens a channel between parties $\mathcal{P}_{1},\mathcal{P}_{2}$. Next, parties exchange with each other signed messages off-chain which update the state of their accounts. If the parties are honest and responsive, they may close the channel in collaboration. To do so, the parties post a mutually signed transaction that spends the output of the funding transaction and awards each party their fair share of funds. However, if a dishonest party $\mathcal{P}_{2}$ attempts to publish on-chain an old state that she profits from comparably to the latest agreed state, her funds will remain locked for the so-called dispute period. During this period, the other (rational) party $\mathcal{P}_{1}$ will try to revoke the old state, by sending a transaction $tx_{1}$ to the mempool called the revocation transaction (or breach remedy). Transaction $tx_{1}$ awards all the channel funds to the cheated party $\mathcal{P}_{1}$. We denote by $f_{1}$ the miner’s fee to include a block with $tx_{1}$.
In this case, the malicious party $\mathcal{P}_{2}$ can launch a timelocked bribing attack, attempting to bribe the miners to ignore $tx_{1}$ for the dispute period $T$ such that $\mathcal{P}_{2}$ gains the additional funds. Specifically, $\mathcal{P}_{2}$ may send in the mempool a block with fee $f_{2}$ that includes transaction $tx_{2}$, with $f_{2}>f_{1}$, that is only valid if no miner includes in their winning block containing $tx_{1}$ within time $T$. Consequently, if the revocation transaction $tx_{1}$ is not published on-chain within $T$, $\mathcal{P}_{2}$ can spend the funds of the old state and the next winning miner will be awarded $f_{2}$. The pair of transactions $tx_{1}$ and $tx_{2}$ is now a conditionally timelocked output.
Assuming that for an average block of transactions, the users get in total $f$ fees, the following holds [10]: if $f_{2}-f>\frac{f_{1}-f}{\lambda_{min}}$ then all rational miners will choose to wait for $T$ rounds and publish a block containing $tx_{2}$.
@@
Feather Fork Attack. A feather fork, as introduced in [24], is an attack on Bitcoin wherein a miner threatens to fork the chain if selected transactions are included. This intimidation mechanism aims to subtly alter the miners’ block acceptance policy: the threatened miners may exclude the selected transactions in order to mitigate the risk of losing their mining reward [34]. As feather forking relies on economic incentives, the attacker may increase their probability of success by bribing other miners to follow their short-lived fork, e.g., committing to pay them the block rewards they may lose by censoring the selected transactions.
@@
Unlike a “hard” fork, where miners exclusively mine their own chain version regardless of its length compared to other versions, a feather fork entails mining on the longest chain that excludes selected transactions and does not fall significantly behind its alternatives. A feather fork is thus less disruptive and more likely to be adopted by the network, making it a potentially powerful tool in the hands of a malicious actor. In this work, we employ this tool to enhance the likelihood of a successful timelocked bribing attack.
2.2 Bribe & Fork Attack
We now introduce our novel attack, termed Bribe & Fork, that combines the timelock bribing attack in LN with the feather fork attack. We assume the existence of a payment channel between parties $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Similarly to timelock bribing, we consider $\mathcal{P}_{2}$ to be malicious and attempt to close the channel with $\mathcal{P}_{1}$ in an old state using the transaction $tx_{2}$. Consequently, $\mathcal{P}_{1}$ is expected to attempt to revoke the old state. To prevent the inclusion of the revocation transaction, $\mathcal{P}_{2}$ bribes the miner $N_{s}$ with the highest mining power $\lambda_{s}$ to threaten others with a fork if they add the unwanted transaction $txs_{1}$ on-chain. This action is implemented through a self-penalty mechanism where the bribed miner temporarily locks collateral which can only be reclaimed in case a block $txs_{2}$ containing $tx_{2}$ appears on-chain (and thus any block $txs_{1}$ containing $tx_{1}$ is not included on-chain). In essence, the bribed miner “bets” that the revocation transaction will be censored, thus rendering the threat credible for the rest of the miners.
To realize Bribe & Fork, there are two mechanisms that should be implementable on-chain: a) the bribe transaction that should only be spendable if $txs_{2}$ is included on-chain, and b) the self-penalty mechanism that enables the bribed miner $N_{s}$ to lock collateral $P$ (with transaction $tx_{p_{1}}$) and then reclaim it only if $txs_{2}$ is included on-chain (with transaction $tx_{p_{2}}$). We implement the bribing and self-penalty mechanisms in Bitcoin script, using the conditioning enabled by the UTXO (Unspent Transaction Output) structure. We note that in Ethereum, preparing a smart contract that has access to the state of the closing channel suffices to implement the bribing and self-penalty mechanisms.
@@ 2.2 Bribe & Fork Attack
In detail, a single bribing transaction $tx_{b}$ and two special transactions $tx_{p_{1}},tx_{p_{2}}$ are introduced. Let us assume that the cheating party $\mathcal{P}_{2}$ is bribing the miners to launch the attack conditioned on the inclusion of its transaction $tx_{2}$. To do so, $\mathcal{P}_{2}$ creates a bribe transaction $tx_{b}$ with input the party’s money from $tx_{2}$ and outputs three UTXOs, one given to the miner that mines this transaction, one dummy output owned by player $N_{s}$ (i.e., the miner bribed to perform feather forking), and one that returns the rest of the money to $\mathcal{P}_{2}$. Now, $N_{s}$ creates a transaction $tx_{p_{1}}$, locking a deposit $P$, that is spendable via a multisignature of $m$-out-of-$n$ parties (e.g., $m=n/2$), one of which is $N_{s}$’s signature. Then, $tx_{p_{2}}$ is created with two inputs: the output of $tx_{p_{1}}$ and the dummy output of $tx_{b}$. Consequently, $tx_{p_{2}}$ is spendable only if it is signed by at least $m$ parties of the predefined set $n$ and $tx_{2}$ is validated on-chain. Upon receiving $tx_{p_{2}}$ signed by at least $m-1$ parties, $N_{s}$ signs and posts it on-chain. Assuming that no subset of size $m-1$ of the rest $n-1$ parties will collude with the miner to spend the deposit, the deposit can be claimed by the miner only if transaction $tx_{2}$ is included on-chain. The security of this scheme depends on the selection of the $n-1$ parties, which can be in principle conditioned on the honest majority assumption of the blockchain via subsampling.
Now, the malicious party $\mathcal{P}_{2}$ together with a selected miner $N_{s}$ can launch the Bribe & Fork attack as follows: They can create special transactions $tx_{b}$ (as a companion transaction for $tx_{2}$) and the self-penalty transactions $tx_{p_{1}},tx_{p_{2}}$ to publicly announce a credible threat that the transaction $tx_{1}$ will be forked once it appears on the blockchain. They publish these special transactions in the mempool as transaction sets $txs_{2}$ (that contains $tx_{2}$ and $tx_{b}$) and $txs_{p_{1}},txs_{p_{2}}$ (that contain $tx_{p_{1}},tx_{p_{2}}$, respectively).
We show later that this attack significantly reduces the cost of the required bribe from $\frac{f_{1}-f}{\lambda_{min}}$ to approximately $\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}$. The required collateral that is eventually reclaimed by the bribed miner is expected to be $\lambda_{s}\cdot B$ (where $B$ is a constant in Bitcoin block reward independent of the user fees), which is comparable to the original bribe needed in [10]. Figure 1 illustrates the comparison of Bribe & Fork with [10], while a figure depicting the details of Bribe & Fork can be found in the Appendix 0.A.

### v2

# Forking Threat Lowers the Bribing Attack Cost

@@ Abstract
In this work, we reexamine the vulnerability of Payment Channel Networks (PCNs) to bribing attacks, where an adversary incentivizes blockchain miners to deliberately ignore a specific transaction to undermine the punishment mechanism of PCNs. While previous studies have posited a prohibitive cost for such attacks, we show that this cost may be dramatically reduced (to approximately \$125), thereby increasing the likelihood of these attacks. To this end, we introduce a modified bribing attack that leverages the threat of a so-called feather fork which we analyze with a novel formal model for the mining game with forking. We empirically analyze historical data of some real-world blockchain implementations to evaluate the scale of this cost reduction. This research sheds more light on the potential vulnerability of PCNs and highlights the need for robust solutions.
@@ 1 Introduction
1.1 Motivation
The financial world was transformed by blockchains such as Bitcoin [nakamoto2008bitcoin] and Ethereum [wood2014ethereum]. While blockchains offer a number of benefits, their scalability remains a significant challenge when compared to traditional centralized payment systems [croman2016scaling]. One promising solution to this issue is the so-called payment channel networks (PCNs) that move most of the transaction workload off-chain [layer2].
A number of PCN proposals have been laid forward so far [aumayr2020generalized, aumayr2021bitcoin, aumayr2021donner, avarikioti2021brick, avarikioti2020cerberus, burchert2018scalable, chakravarty2020hydra, decker2019eltoo, decker2015fast, dong2018celer, dziembowski2019perun, dziembowski2018general, egger2019atomic, Herlihy:2018:ACC:3212734.3212755, jourenko2020lightweight, poon2016bitcoin, spilman2013anti], each design offering some unique combination of features. Nonetheless, the core idea behind payment channels remains the same, that is to facilitate off-chain transactions among parties connected, either directly or indirectly via a path, on a network operating on top of the blockchain layer, the PCN.
@@ 1.1 Motivation
Hence, the security of PCNs, like the most widely-deployed Bitcoin Lighting Network [poon2016bitcoin], crucially relies on financial incentives. Specifically, during the dispute period, the punishment mechanism should enforce that a malicious party is always penalized and an honest party should never lose its funds. Unfortunately, this is not always the case, as argued by Khabbazian et al. [timelocked] and Avarikioti et al. [zeta]. For instance, Lighting channels are susceptible to the so-called timelock bribing attacks. In such an attack, a malicious party posts an old update transaction on-chain, attempting to close its channel with more funds than it presently possesses. Concurrently, the party bribes the miners to ignore the punishment transaction of the counterparty for the duration of the dispute period. This bribe is typically offered to the miners in the form of high transaction fees.
Naturally, the success of this attack depends on the value of the bribe. Avarikioti et al. [zeta] showed that a bribing attack will be successful if the bribe is no smaller than:
$\frac{f_{1}-f}{\lambda_{min}},$
where $\lambda_{min}$ is the fraction of the mining power controlled by the least significant miner in the underlying proof-of-work blockchain, $f_{1}$ is the sum of the fees of a block containing the punishment transaction, and $f$ is the sum of the fees of a block containing average transactions.
@@
This result implies that the bribing attack is unlikely to happen in practice. That is because the mining power $\lambda_{min}$ of the least significant miner may be arbitrarily small, leading to bribing amounts that may significantly exceed the funds typically locked in PCNs. But even in blockchains where the mining power is rather concentrated like in Bitcoin [long2022measuring], the cost of a bribing attack is very high. For example, if we very conservatively assume that the smallest miner has $10^{-4}$ of the total mining power [^1] In the Section 5.2, we experimentally show, that the mining power of the weakest miner in the system can be fairly assumed to be of magnitude $10^{-12}$., the cost of the attack as analyzed in [zeta] would be at least $1$ BTC, for $f_{1}-f\approx 10^{-4}$ BTC. As a result, it would be irrational to perform a bribing attack of this sort, as the average closing price for $1$ BTC between, for instance, 2019 and 2022 was $23,530.92$ USD222statmuse.com, which is more than 10-fold the current total value locked on average in a Lighting channel [^3] https://1ml.com/statistics. This naturally leads to questioning whether there is potential to amplify such attacks to the extent they pose a genuine threat to the security of PCNs.
1.2 Our Contribution
[…]
@@ 1.2 Our Contribution
where $\overline{f}$ is the average transaction fee, and $\lambda_{s}$ is the mining fraction of the most significant miner. Recall that $f_{1}$ and $f$ denote the sum of fees of a block containing the punishment transaction and only average transactions respectively. To demonstrate this improvement with an example, let us consider Bitcoin to fairly compare to the numbers presented above. We estimate $\overline{f}$ is approximately $10^{-4}$ BTC, while $\lambda_{s}$ is approximately $20\%$. $\lambda_{s}$ replaces in the denominator the previously presented $\lambda_{min}<<10^{-4}$, thus yielding a bribe at least 1000 times smaller than the one in [zeta].
To derive this result, we present a formal model of mining games with forking, extending the conditionally timelocked games introduced by Avarikioti et al. [zeta]. In the game with forks, miners may now choose, in each round, (a) which transactions to mine, and (b) whether they want to continue one of the existing chains or they intend to fork one of the chains. All miners are aware of the choices of the winner of each block, as a feather-forking miner locks collateral on-chain.
To empirically estimate the cost reduction of our novel bribing attack, we analyze the historical data of real-world blockchain implementations. Among others, we analyze the average block rewards and fees, as well as the hash power present in the system and available to a single miner, primarily for Bitcoin in 2022. Given the officially available data, we observe that the cost of our attack can be as cheap as \$125 (for 1 BTC $\approx$ \$25.000). In general, the cost of our attack can up to $10^{10}$ times cheaper than the bribe required in [zeta] according to our findings. Therefore, even when considering the collateral being around \$$\$30,000$, our bribing attack is substantially more cost-efficient, and by extension, more probable to occur.
2 Background and Assumptions
In this section, we provide the necessary background and assumptions. We first describe proof-of-work (PoW) mining, the timelocked bribing attack, and the feather fork attack (Section 2.1). Then, we provide a brief description of the Bitcoin Lightning Network and how timelocked bribing attacks affect its security according to [zeta, timelocked] (Section 2.2). We conclude with our key assumptions that are necessary to systematically model and study the behavior of players in our proposed bribing attack (Section 2.4).
2.1 PoW Mining and Attacks
We solely focus on proof-of-work blockchains that employ the so-called Nakamoto consensus extending the longest chain, such as Bitcoin [nakamoto2008bitcoin]. We assume that in such environments miners tend to form mining pools to (a) bypass the task of verifying transactions – the pool’s manager dispatches a list of valid transactions for inclusion – and more importantly, (b) to guarantee a more stable income as individual mining carries substantial deviation. Empirical evidence supporting this assumption, drawn from Bitcoin, will be detailed in Section 5. For the rest of this work, each miner is assumed to either mine independently or stick to a selected mining pool for the whole course of the game, i.e., we will treat the mining pools as single players in the game. We will use players and miners interchangeably.
@@ 2.1 PoW Mining and Attacks
Feather Fork Attack. A feather fork, originally introduced in [forking_discussion], is an attack on the Bitcoin network wherein a miner threatens to fork the chain if selected transactions are included. This intimidation mechanism aims to subtly alter the miners’ block acceptance policy: the threatened miners may exclude the selected transactions in order to mitigate the risk of losing their mining reward [temporary]. As feather forking relies on economic incentives, the attacker may increase their probability of success by bribing other miners to follow their short-lived fork, e.g., committing to pay them the block rewards they may lose by censoring the selected transactions.
@@
Unlike a “hard” fork, where miners exclusively mine their own chain version regardless of its length compared to other versions, a feather fork entails mining on the longest chain that excludes selected transactions and is not falling significantly behind its alternatives. A feather fork is thus less disruptive and more likely to be adopted by the network, making it a potentially powerful tool in the hands of a malicious actor. In this work, we employ this tool to enhance the likelihood of a successful timelocked bribing attack.
2.2 The Bitcoin Lightning Network
Payment Channel Networks (PCN) are protocols deployed on top of blockchains to increase their transaction throughput. In this work, we focus on PCN instantiations that operate like the most widely-deployed PCN, namely the Bitcoin Lightning Network (LN).
In LN, a single on-chain transaction called the funding transaction, opens a channel between parties $\mathcal{P}_{1},\mathcal{P}_{2}$. Next, parties exchange with each other signed messages off-chain which update the state of their accounts. If the parties are honest and responsive, they may close the channel in collaboration. To do so, the parties post a mutually signed transaction that spends the output of the funding transaction and awards each party their fair share of funds. However, if a dishonest party $\mathcal{P}_{2}$ attempts to publish on-chain an old state that she profits from comparably to the latest agreed state, her funds will remain locked for the so-called dispute period. During this period, the other (rational) party $\mathcal{P}_{1}$ will try to revoke the old state, by sending a transaction $tx_{1}$ to the mempool called the revocation transaction (or breach remedy). Transaction $tx_{1}$ awards all the funds of the channel to the cheated party $\mathcal{P}_{1}$. We assume the fee of the miner for including $tx_{1}$ is $f_{1}$.
In this case, the malicious party $\mathcal{P}_{2}$ can launch a timelocked bribing attack, attempting to bribe the miners to ignore $tx_{1}$ for the dispute period $T$ such that $\mathcal{P}_{2}$ gains the additional funds. Specifically, $\mathcal{P}_{2}$ may send in the mempool a transaction $tx_{2}$ with fee $f_{2}$, with $f_{2}>f_{1}$, that is only valid if no miner includes in their winning block $tx_{1}$ within time $T$. Consequently, if the revocation transaction $tx_{1}$ is not published on-chain within time $T$, the malicious party can spend the funds of the old state and the next winning miner will be awarded the fee $f_{2}$. The pair of transactions $tx_{1}$ and $tx_{2}$ is now a conditionally timelocked output.
Assuming that for an average block of transactions, the users get in total $f$ fees, according to Avarikioti et al. [zeta] the following holds: if $f_{2}-f>\frac{(f_{1}-f)}{\lambda_{min}}$ then all rational miners will choose to wait for $T$ rounds and publish a block containing $tx_{2}$.
2.3 Bribing and Self-Penalty Mechanism
The work of Avarikioti et al. [zeta] introduced a special bribing transaction that allows to control how the dishonest funds are spent. In our analysis, we additionally assume that the miner with the biggest mining power (we call him the player $N_{s}$ in this paragraph) can create a pair of special (sets of) transactions $(tx_{p_{1}},tx_{p_{2}})$ implementing the self-penalty mechanism. This mechanism assumes a pair of conditionally timelocked transactions $(tx_{1},tx_{2})$ and a parameter $P$. The transaction $tx_{p_{1}}$ locks a penalty $P$ of the player’s $N_{s}$ funds. The pair of special transactions $(tx_{p_{1}},tx_{p_{2}})$ has a property that the player $N_{s}$ is able to withdraw the locked funds using the transaction $tx_{p_{2}}$ only after the transaction $tx_{2}$ appears on the blockchain.
@@ 2.3 Bribing and Self-Penalty Mechanism
To implement the bribing and self-penalty mechanism in Bitcoin script, we use the conditioning enabled by the UTXO (Unspent Transaction Output) structure. In detail, a single bribing transaction $tx_{b}$ and two special transactions $tx_{p_{1}},tx_{p_{2}}$ are introduced. Let us assume that the cheating party $\mathcal{P}_{2}$ is bribing the miners to launch the attack conditioned on the inclusion of its transaction $tx_{2}$. To do so $\mathcal{P}_{2}$ creates a bribe transaction $tx_{b}$ with input the party’s money from $tx_{2}$ and output three UTXOs, one given to the miner that mines this transaction, one dummy output owned by player $N_{s}$ (i.e., the miner bribed to perform feather forking), and one that returns the rest of the money to $\mathcal{P}_{2}$. Now, $N_{s}$ creates a transaction $tx_{p_{1}}$, locking a deposit $P$, that is spendable via a mutisignature of $m$-out-of-$n$ parties (e.g., $m=n/2$), one of which is $N_{s}$’s signature. Then, $tx_{p_{2}}$ is created with two inputs: the output of $tx_{p_{1}}$ and the dummy output of $tx_{b}$. Consequently, $tx_{p_{2}}$ is spendable only if it is signed by at least $m$ parties of the predefined set $n$ and $tx_{2}$ is validated on-chain. Upon receiving $tx_{p_{2}}$ signed by at least $m-1$ parties, $N_{s}$ signs and posts it on-chain. Assuming that no subset of size $m-1$ of the rest $n-1$ parties will collude with the miner to spend the deposit, the deposit can be claimed by the miner only if transaction $tx_{2}$ is included on-chain. The security of this scheme depends on the selection of the $n-1$ parties, which can be in principle conditioned on the honest majority assumption of the blockchain via subsampling.
We note that in Ethereum, preparing a smart contract that has access to the state of the closing channel suffices to implement the penalty mechanism.
2.4 Key Assumptions
As we model the complex process of PoW blockchain mining, we make a few assumptions that let us systematically study behavior of the players who take part in this process.
Average transactions. In Section 5.1, based on empirical data, we show how block rewards fluctuate in the real-world. However, following [zeta], we assume that standard transactions have a constant (average) reward and that all blocks have a constant number of transactions. We refer to each such standard transaction as an average transaction.
Abandon rules for forked states. In general, miners have the option to fork a chain or mine one of the existing chains from an arbitrary block of the chain. However, forking that starts at some old block is unrealistic; hence, we exclude this possibility from the game (following [forking_discussion, temporary, forkanalysis]). In other words, miners in our model can fork only the last transaction on one of the chains, and then either the original chain or the fork becomes stable whenever it reaches a length equal to the length of the original chain plus one. One can also make an attempt to model forks differently, e.g. with a finite automaton definition. We expect that our results also hold for the models with longer forks or models based on finite automaton definition, although the parameters of our solution could change.
Mining structure. As already mentioned in the previous section, we assume in this work that each miner either mines independently or sticks to a selected mining pool for the whole course of the game. As conducting a feather-fork attack requires a credible threat, it should be conducted by a player with relatively strong mining power. From our empirical analysis of real-life cryptocurrencies, we estimate that typically there exists a miner that has a mining power of at least $20\%$ of the total hash power. Next, we assume that there exists at least one other player in the game that has a mining power slightly higher than $1\%$ and that no players with overwhelmingly small mining power (less than a $10^{-100}$ fraction of the total mining power) participate in the game. In Section 5.2, we show a detailed analysis of the distribution of miners’ computing power on real-life data.
Preventing shallow forks. In Conjecture 1, we assume that players will have the option to fork some transaction only when they see an explicit opportunity of mining some transaction in the future, initially blocked by the currently forked transaction. In other words, we assume that players will not start fork what we call “average” transactions. That is, we forbid shallow forks in the model. This is because, in the real-world, excessive forking can trigger a denial of service attack on the blockchain. Thus, it is very unlikely that miners will follow the strategy of forking arbitrary transactions.
2.5 Bribe & Fork Overview
Missing text